# Towards Model Extraction Attacks in GAN-Based Image Translation via Domain Shift Mitigation

### 19 Mar 2024 v3

@@ 1 Introduction
To close the distribution gap between query and training samples, numerous works have been proposed for various classification tasks, including image processing (Orekondy, Schiele, and Fritz 2019; Pal et al. 2020; Yuan et al. 2022; Barbalau et al. 2020), NLP processing (Krishna et al. 2019; Xu et al. 2021), and graph data processing (Wu et al. 2022; Shen et al. 2022). For instance, in image classification, (Orekondy, Schiele, and Fritz 2019) suggested using reinforcement learning, and (Pal et al. 2020) suggested using active learning as a means to identify better query samples. Alternatively, (Yuan et al. 2022; Barbalau et al. 2020) advocated aligning the distribution of the secret training dataset with Generative Adversarial Networks (GANs). In NLP tasks, (Xu et al. 2021) suggested targeting an ensemble of victim APIs simultaneously, all providing the same service, to make them implicitly vote for good query samples.
Concurrent with classification tasks, image-to-image translation (I2IT) tasks, such as image super-resolution or restoration, constitute another significant application domain of DNNs. As a testament to this prevalence, Table 2 makes a comparison of the pay-per-use price between classification and I2IT tasks on different platforms [^1] Providing image translation service: https://cloud.baidu.com, https://imglarger.com, ${}^{,}$ [^2] Providing classification service: https://cloud.baidu.com, https://aws.amazon.com, https://cloud.google.com.. Nonetheless, the MEA vulnerabilities in I2IT are rarely explored, making their risk largely underestimated.
@@
Table 1: Pay-per-use price (/image) comparison on different APIs11footnotemark: ${}^{,}$22footnotemark: 2.
@@ 2 Background and Related Work
$\mathcal{L}_{\text{GAN}}=\min_{G}\max_{D}~{}$ ${\mathbb{E}}_{x\sim P_{\text{data}}(x)}[{\rm log}D(x)]$ (1)
@@
As mentioned in Sec. 1, by taking the victim model $F_{V}$ as a labeling oracle, the goal of MEA involves creating an attack model $F_{\mathcal{A}}$ that emulates the functionalities of the victim model $F_{V}$. In classification tasks, this goal translates to finding the decision boundary while reducing the query budget under the assumption of attackers’ knowledge. Strategies include selecting better query samples with reinforcement learning (Orekondy, Schiele, and Fritz 2019) or active learning (Pal et al. 2020), or even harnessing GAN to generate images that are close to the secret training data of $F_{V}$ (Yuan et al. 2022; Barbalau et al. 2020).
@@ 4 Method
The domain shift problem, as discussed in (Glorot, Bordes, and Bengio 2011), typically pertains to the decline in model performance caused by disparities between the distribution of training data and that of test data. In this work, we identify this phenomenon as a fundamental factor affecting the performance of MEA. In MEAs, attackers lack access to the secret data used to train the victim model and they usually only use public datasets as attack data. When a disparity exists between the distribution of this public data and the distribution of the secret data, it causes the attack model’s output quality to decrease and can even result in deviations from the target domain. Such data that fails to effectively represent the mapping relationship between the source and target domains can be considered noise data, which hinders the training of subsequent attack models.
@@
For MEA of classification tasks, traditional wisdom aims at closing the distributional gap between the victim dataset and the attack dataset (Orekondy, Schiele, and Fritz 2019; Pal et al. 2020; Yuan et al. 2022). Due to the unique challenge of I2IT tasks, this work aims to address the impact of the domain shift problem from an orthogonal perspective by resorting to a flatter and smoother loss landscape for the attack model. Fig. 2 illustrates our insight. In the sketch, we briefly illustrate the effect of different factors on the model’s test loss. Under vanilla training of the attack model, the obtained $F_{\mathcal{A}}$ will be fitted to the distribution of the attack’s training set $D_{\mathcal{A}}$, hence deviating from the original optima of $F_{V}$. In contrast, when the model is enforced to be with lower complexity (i.e., smooth) around a flat minimum area, such local overfitting introduced by domain shift can be mitigated.
@@
We then train the attack model using sharpness-aware minimization (SAM). Due to the noise caused by domain shifting (i.e., $D_{V}\rightarrow D_{\mathcal{A}}$), the inherent issue of mode collapse during GAN training can be exacerbated, which leads the model to overly specialize in certain patterns during the generation process, resulting in an overfitting to partial data (d’Ascoli, Sagun, and Biroli 2020). Recent studies have demonstrated that SAM (Foret et al. 2020) optimizer has held the promise of seeking out flatter minima by simultaneously minimizing loss value and loss sharpness. However, to the best of our knowledge, there have been no prior works reporting the process of training GANs using SAM. We therefore introduce a GAN-specific SAM variant towards a wide minimum and further improve the effectiveness of MEA against the I2IT network.
@@
[…]
@@
[…]
@@ 5 Experimental Analyses
To build the victim model, we use CycleGAN (Liu, Breuel, and Kautz 2017) to train the style transfer tasks, and we directly employ the pre-trained real-ESRGAN model (Wang et al. 2021) for the super-resolution task. The datasets employed to train the victim style transfer models are horse2zebra and photo2vangogh in (Zhu and et al. 2021). The datasets used for training the victim super-resolution model are DIV2K (Agustsson and Timofte 2017), Flickr2K (Timofte and et al. 2017), and OutdoorSceneTraining (Wang et al. 2018b). All of these are identical to the original settings (Liu, Breuel, and Kautz 2017; Wang et al. 2021). All victim models have demonstrated compelling performance in their respective tasks.
To train the model, we use Pix2Pix (Isola et al. 2017) and CycleGAN (Liu, Breuel, and Kautz 2017) as attack modeling backbone frameworks for style transfer and super-resolution tasks, respectively.
@@
Table 2: Quantitative results of our attack against the style transfer $\ddagger$ (A-C represents Artist-Copy).
@@ 6 MEA Agaisnt I2IT in Real Life
We randomly select 2,000 images from the face dataset FFHQ (512$\times$512) (Marinez 2020) and Landscape dataset (Rougetet 2020) as the attack dataset, respectively. The face dataset FFHQ is composed of 52,000 high-quality PNG images with 512$\times$512 resolution crawled from Flickr. These images vary considerably in age, race, and image background, and are automatically aligned and cropped using dlib. We use the images in this dataset as the attack dataset for both face2cartoon (2d) and face2cartoon (3d) tasks and choose CycleGAN as the attack model backbone. The Landscape dataset is the same we used in the photo2vangog task, and Pix2Pix is used as the attack model backbone for landscape2cartoon task. To evaluate the extraction performance, we take 200 and 150 images from the remaining parts of the FFHQ and landscape datasets, respectively, as the test set.

### 16 Mar 2026 v4

@@ 1 Introduction
To close the distribution gap between query and training samples, numerous works have been proposed for various classification tasks, including image processing (Orekondy et al. 2019; Pal et al. 2020; Yuan et al. 2022; Barbalau et al. 2020), NLP processing (Krishna et al. 2019; Xu et al. 2021), and graph data processing (Wu et al. 2022; Shen et al. 2022). For instance, in image classification, (Orekondy et al. 2019) suggested using reinforcement learning, and (Pal et al. 2020) suggested using active learning as a means to identify better query samples. Alternatively, (Yuan et al. 2022; Barbalau et al. 2020) advocated aligning the distribution of the secret training dataset with Generative Adversarial Networks (GANs). In NLP tasks, (Xu et al. 2021) suggested targeting an ensemble of victim APIs simultaneously, all providing the same service, to make them implicitly vote for good query samples.
Concurrent with classification tasks, image-to-image translation (I2IT) tasks, such as image super-resolution or restoration, constitute another significant application domain of DNNs. As a testament to this prevalence, Table 2 makes a comparison of the pay-per-use price between classification and I2IT tasks on different platforms [^1] Providing image translation service: https://cloud.baidu.com, https://imglarger.com, https://vanceai.com., [^2] Providing classification service: https://cloud.baidu.com, https://aws.amazon.com, https://cloud.google.com.. Nonetheless, the MEA vulnerabilities in I2IT are rarely explored, making their risk largely underestimated.
@@
Table 1: Pay-per-use price (/image) comparison on different APIs11footnotemark: 1,22footnotemark: 2.
@@ 2 Background and Related Work
$\mathcal{L}_{\text{GAN}}=\min_{G}\max_{D}$ ${\mathbb{E}}_{x\sim P_{\text{data}}(x)}[{\rm log}D(x)]$ (1)
@@
As mentioned in Sec. 1, by taking the victim model $F_{V}$ as a labeling oracle, the goal of MEA involves creating an attack model $F_{\mathcal{A}}$ that emulates the functionalities of the victim model $F_{V}$. In classification tasks, this goal translates to finding the decision boundary while reducing the query budget under the assumption of attackers’ knowledge. Strategies include selecting better query samples with reinforcement learning (Orekondy et al. 2019) or active learning (Pal et al. 2020), or even harnessing GAN to generate images that are close to the secret training data of $F_{V}$ (Yuan et al. 2022; Barbalau et al. 2020).
@@ 4 Method
The domain shift problem, as discussed in (Glorot et al. 2011), typically pertains to the decline in model performance caused by disparities between the distribution of training data and that of test data. In this work, we identify this phenomenon as a fundamental factor affecting the performance of MEA. In MEAs, attackers lack access to the secret data used to train the victim model and they usually only use public datasets as attack data. When a disparity exists between the distribution of this public data and the distribution of the secret data, it causes the attack model’s output quality to decrease and can even result in deviations from the target domain. Such data that fails to effectively represent the mapping relationship between the source and target domains can be considered noise data, which hinders the training of subsequent attack models.
@@
For MEA of classification tasks, traditional wisdom aims at closing the distributional gap between the victim dataset and the attack dataset (Orekondy et al. 2019; Pal et al. 2020; Yuan et al. 2022). Due to the unique challenge of I2IT tasks, this work aims to address the impact of the domain shift problem from an orthogonal perspective by resorting to a flatter and smoother loss landscape for the attack model. Fig. 2 illustrates our insight. In the sketch, we briefly illustrate the effect of different factors on the model’s test loss. Under vanilla training of the attack model, the obtained $F_{\mathcal{A}}$ will be fitted to the distribution of the attack’s training set $D_{\mathcal{A}}$, hence deviating from the original optima of $F_{V}$. In contrast, when the model is enforced to be with lower complexity (i.e., smooth) around a flat minimum area, such local overfitting introduced by domain shift can be mitigated.
@@
We then train the attack model using sharpness-aware minimization (SAM). Due to the noise caused by domain shifting (i.e., $D_{V}\rightarrow D_{\mathcal{A}}$), the inherent issue of mode collapse during GAN training can be exacerbated, which leads the model to overly specialize in certain patterns during the generation process, resulting in an overfitting to partial data (d’Ascoli et al. 2020). Recent studies have demonstrated that SAM (Foret et al. 2020) optimizer has held the promise of seeking out flatter minima by simultaneously minimizing loss value and loss sharpness. However, to the best of our knowledge, there have been no prior works reporting the process of training GANs using SAM. We therefore introduce a GAN-specific SAM variant towards a wide minimum and further improve the effectiveness of MEA against the I2IT network.
@@
$\left\{\begin{aligned} &\epsilon(\mathbf{w}^{G_{i}})=\rho_{G_{i}}\cdot\frac{\nabla_{\mathbf{w}^{G_{i}}}L(\mathbf{w})}{\|\nabla_{\mathbf{w}^{G_{i}}}L(\mathbf{w})\|_{2}},\\ &g(\mathbf{w}^{G_{i}})=\nabla_{\mathbf{w}^{G_{i}}}L(\mathbf{w}^{G_{i}})|_{\mathbf{w}^{G_{i}}+\epsilon(\mathbf{w}^{G_{i}})},\\ &\mathbf{w}_{t+1}^{G_{i}}=\mathbf{w}_{t}^{G_{i}}-\alpha_{t}\cdot g(\mathbf{w}^{G_{i}}),\end{aligned}\right$.
@@
$\left\{\begin{aligned} &\epsilon(\mathbf{w}^{D_{j}})=\rho_{D_{j}}\cdot\frac{\nabla_{\mathbf{w}^{D_{j}}}L(\mathbf{w})}{\|\nabla_{\mathbf{w}^{D_{j}}}L(\mathbf{w})\|_{2}},\\ &g(\mathbf{w}^{D_{j}})=\nabla_{\mathbf{w}^{D_{j}}}L(\mathbf{w}^{D_{j}})|_{\mathbf{w}^{D_{j}}+\epsilon(\mathbf{w}^{D_{j}})},\\ &\mathbf{w}_{t+1}^{D_{j}}=\mathbf{w}_{t}^{D_{j}}+\alpha_{t}\cdot g(\mathbf{w}^{D_{j}}),\end{aligned}\right$.
@@ 5 Experimental Analyses
To build the victim model, we use CycleGAN (Liu et al. 2017) to train the style transfer tasks, and we directly employ the pre-trained real-ESRGAN model (Wang et al. 2021) for the super-resolution task. The datasets employed to train the victim style transfer models are horse2zebra and photo2vangogh in (Zhu and et al. 2021). The datasets used for training the victim super-resolution model are DIV2K (Agustsson and Timofte 2017), Flickr2K (Timofte and et al. 2017), and OutdoorSceneTraining (Wang et al. 2018b). All of these are identical to the original settings (Liu et al. 2017; Wang et al. 2021). All victim models have demonstrated compelling performance in their respective tasks.
To train the model, we use Pix2Pix (Isola et al. 2017) and CycleGAN (Liu et al. 2017) as attack modeling backbone frameworks for style transfer and super-resolution tasks, respectively.
@@
[…]
@@ 6 MEA Agaisnt I2IT in Real Life
We randomly select 2,000 images from the face dataset FFHQ ($\times512$) (Marinez 2020) and Landscape dataset (Rougetet 2020) as the attack dataset, respectively. The face dataset FFHQ is composed of 52,000 high-quality PNG images with $\times512$ resolution crawled from Flickr. These images vary considerably in age, race, and image background, and are automatically aligned and cropped using dlib. We use the images in this dataset as the attack dataset for both face2cartoon (2d) and face2cartoon (3d) tasks and choose CycleGAN as the attack model backbone. The Landscape dataset is the same we used in the photo2vangog task, and Pix2Pix is used as the attack model backbone for landscape2cartoon task. To evaluate the extraction performance, we take 200 and 150 images from the remaining parts of the FFHQ and landscape datasets, respectively, as the test set.